# Improving Underwater Acoustic Classification Through Learnable Gabor Filter Convolution and Attention Mechanisms

### v1

@@ Abstract
Remotely detecting and classifying underwater acoustic targets is critical for environmental monitoring and defence. However, the complex nature of ship-radiated and environmental underwater noise poses significant challenges to accurate signal processing. While recent advancements in machine learning have improved classification accuracy, issues such as limited dataset availability and a lack of standardised experimentation hinder generalisation and robustness. This paper introduces GSE ResNeXt, a deep learning architecture integrating learnable Gabor convolutional layers with a ResNeXt backbone enhanced by squeeze-and-excitation attention mechanisms. The Gabor filters serve as two-dimensional adaptive band-pass filters, extending the feature channel representation. Its combination with channel attention improves training stability and convergence while enhancing the model’s ability to extract discriminative features. The model is evaluated on three classification tasks of increasing complexity. In particular, the impact of temporal differences between the training and testing data is explored, revealing that the distance between the vessel and sensor significantly affects performance. Results show that, GSE ResNeXt consistently outperforms baseline models like Xception, ResNet, and MobileNetV2, in terms of classification performance. Regarding stability and convergence, the addition of Gabor convolutions in the initial layers of the model represents a 28% reduction in training time. These results emphasise the importance of signal processing strategies in improving the reliability and generalisation of models under different environmental conditions, especially in data-limited underwater acoustic classification scenarios. Future developments should focus on mitigating the impact of environmental factors on input signals.
@@ 1 Introduction
Remote detection of underwater acoustic targets is a valuable and fundamental asset for environmental protection and defence, as its applications include surveillance, dredging, hydrography, marine life research, and fisheries protection, to name a few (bjorno2017chapter). Given the sensitive nature of these applications, the correct detection, processing, and classification of such signals is essential. In the defence field, for instance, incorrect classification could lead to false alarms or even to the non-identification of a potential threat. In environmental protection contexts, misclassifications can directly affect species population estimates, possibly leading to inappropriate management strategies (caillat2013effects).
@@
Recently, a significant increase in anthropogenic activities conducted in the ocean has elevated the levels of underwater acoustic noise (ross2005ship; hildebrand2009anthropogenic), driving new studies to focus on the control of maritime traffic by detecting and classifying ship types by their radiated noise. The complexity of such signals, on the other hand, poses a challenge to their effective management in certain situations. Noise recorded by passive sonars is usually composed of environmental sounds and ship-radiated noise (domingos2022survey). Environmental sounds include sea surface noise, such as breaking waves and wind, and intermittent biological sounds, which depend on factors such as location, weather conditions, and seasonality (xu2016digital). Ship-emitted noise, on the other hand, consists of machinery and propeller noises, which appear as constant lower components in the frequency spectra, and hydrodynamic noise, which is a broadband component that shifts between mid and high frequencies according to the ship’s speed (carlton2018marine). Additionally, multi-path sound reflection caused by interaction with different media, such as the sea surface, can affect classification accuracy within recordings (wilmut2007inversion; mckenna2012underwater).
@@
With the evolution of machine learning methods, their application in underwater target classification has been increasing. With a focus on feature extraction and classification of ship types, recent research has shown an interest in improving the machine learning architecture to extract relevant information from complex input data. A subset of works uses autoencoders to improve the feature extraction process (luo2020underwater; li2020feature; irfan2021deepship), reducing the need for manual hyperparameters tuning. These architectures take advantage of the data-driven training process, which automatically extracts features from the input data, minimising a given loss function. However, the acquisition of such signals is sometimes sensitive and complex, leading to limited available data and, therefore, limited features. Another common strategy applied for detection and classification is the use of recurrent neural network blocks to extract the temporal dependency of the acoustic signal (liu2021underwater; alouani2022spatiotemporal; han2022underwater). This strategy aims to extract features based on the temporal components of the audio; however, as the ship noise is almost constant over short periods of time, the improvements by focusing on short-time dependencies are minimal.
@@
Notably, the most common approach to detecting and classifying underwater acoustic signals involves an initial pre-processing stage, where the raw waveform is transformed into a spectrogram. The resulting two-dimensional representation allows the signal to be classified using algorithms originally developed for computer vision tasks. Residual neural networks (ResNet) (he2016deep) serve as the baseline for several pipelines in underwater acoustic target classification (hong2021underwater; yao2023underwater; ren2022ualf; domingos2022investigation). They introduce identity shortcut connections between convolutional layers, enabling the network to bypass one or more layers by directly passing the output of earlier layers to later ones. These connections extend the model’s capacity by allowing the exploration of alternative paths during inference, thereby mitigating the vanishing gradient problem (veit2016residual). Subsequent improvements introduced the concept of cardinality (xie2017aggregated), which enhances the model’s ability to explore multiple learning paths through grouped convolution blocks within residual layers. The resulting model, ResNeXt, often achieves higher classification accuracy with only a modest increase in computational cost. In underwater acoustics, the application of ResNets has proven highly beneficial for improving classification accuracy, as they enable deeper architectures and facilitate the exploration of multiple inference paths (hong2021underwater; yao2023underwater; ren2022ualf; domingos2022investigation). In a similar fashion, multi-branch classification approaches are also frequently employed (xie2024unraveling; xie2024advancing; wang2023underwater; tian2023joint), further enhancing the multi-path capability of the pipelines. The successful application of these strategies highlights the feasibility of employing computer vision techniques for classification tasks based on spectrograms.
@@
Attention mechanisms have been incorporated into classification pipelines, allowing models to dynamically assign importance to input elements and focus on contextually relevant information when producing outputs (bahdanau2016neural). These mechanisms have been further emphasised with the advent of transformer models (vaswani2017attention), which rely solely on attention mechanisms and dispense with convolutions. Despite the impressive results obtained in vision applications using transformers, the combination of attention mechanisms and convolutional neural networks is still being explored, yielding superior results in several tasks (han2023survey). In vision tasks, attention is typically applied along the spatial dimension, assigning adaptive importance factors across the image (zhu2019empirical; dosovitskiy2021image), or along the channel dimension, prioritising different feature channels within the neural network (wang2020ecanet; hu2018squeezeandexcitation). Subsequent approaches also focus on combining both strategies in a spatial–channel framework (woo2018cbam; chen2017scacnn), which aims to enhance the model’s capability to represent and extract features from the inputs. For spectrograms, as the spatial dimension has different scales along the vertical and horizontal axes, some applications separate the time and frequency dimensions when applying spatial attention methods (yang2024underwater). In underwater acoustic classification, attention mechanisms have proven effective in improving accuracy by assigning different levels of importance to features at local and global scales (zhou2023attentionbased; yang2023lightweight; xu2023underwater; yang2024underwater). While these techniques continue to advance classification performance, their effectiveness is highly dependent on the type of spectral representation used (domingos2022investigation). Processing these representations to address the complex behaviour of underwater acoustic signals offers an attractive alternative to improvements in neural networks alone.
@@
Gabor filters, which are band-pass filters well suited to local frequency and orientation analysis (daugman1985uncertainty), have the potential to enhance feature extraction in convolutional neural networks. Their application in underwater acoustic tasks has also been explored in recent years. In a study of whale vocalisations, Gabor filtering was applied to enhance distributed acoustic sensing (DAS) data, improving the signal-to-noise ratios of the inputs (goestchel2025enhancing). While it supports the broader applicability of these filters in underwater environments, DAS data differs fundamentally from conventional hydrophone data, which was not addressed by this work. In a hydrophone-based classification problem, Gabor filters were applied directly to raw audio as a preprocessing step, generating a single-channel spectral representation as input (elsborg2025acoustic). This preprocessing stage dynamically emphasised different frequency components, improving classification performance. However, the resulting spectral input was used without further processing, leaving the task of feature extraction entirely to the model. The application of these filters has also been studied in the two-dimensional domain, where they were applied to spectrogram inputs followed by a low-pass filter (ren2022ualf). Despite achieving a 3.5% improvement in accuracy, this approach focused solely on enhancing the initial representations, without further processing the resulting frequency bands. These results support the use of Gabor filters to improve generalisation, but leave open the question of how the different frequency bands can be more effectively exploited within the neural network architecture. Furthermore, most previous works consider only a single classification task, which limits the generalisation of their findings.
@@
This paper presents a method for underwater acoustic classification that optimises the feature extraction layer, providing an experimentation strategy that addresses three different scenarios and considers the variability of the training dataset. The Gabor Squeeze-and-Excitation ResNeXt model (GSE ResNeXt) introduces the usage of Gabor convolutional filters in the initial layers of the neural network combined with residual layers and channel attention mechanisms in the models’ backbone, showing to be effective at improving the training stability and classification performance. Additionally, the use of Gabor filters facilitated the convergence of the model, representing a substantial enhancement for future applications that utilise limited number of training examples. In summary, the contributions of this paper are as follows:
@@
    The application of learnable Gabor convolutional layers into the underwater acoustic classification problem, improving the convergence and stability of the model during the training process.
@@
    The incorporation of channel attention mechanisms alongside Gabor convolutional layers, facilitating the feature extraction process by serving as a two-dimensional frequency selector, thereby improving classification performance.
@@
    A robust analysis of the classification results, focusing on the temporal consistency of the predictions across three tasks of increasing complexity using real-world data.
@@ 2.1 Two-dimensional Gabor convolution
Traditional convolutional blocks perform the convolution of a learnable two-dimensional kernel on an input image, occasionally adding an additional bias term depending on the application (lecun2015deep). When visualising the convergence of convolutional networks, it was observed that the convolutional kernels in the initial layers of models trained for image classification converged to Gabor-like filters and colour blob representations (krizhevsky2012imagenet; zeiler2013visualizing). These findings suggest that the training process could be enhanced by constraining the spatial kernels to Gabor functions, which possess significantly fewer parameters to optimise and are capable of extracting spatial textures from visual inputs. Building on these conclusions, although several studies have employed Gabor functions to pre-process and extract features from images, the integration of a Gabor filter layer with learnable weights into the back-propagation algorithm was only proposed years later (alekseev2019gabornet). Applying this approach to image classification tasks demonstrated improved robustness, as the visual kernels were less sensitive to noisy information, while also reducing training complexity.
A Gabor function can be understood as a complex sinusoid modulated by a Gaussian envelope. While the complex sinusoid captures the frequency and phase information of the features, the Gaussian component localises them in space. In image processing, the imaginary part of the sinusoid is often omitted to simplify the calculation process, since it is sufficient for texture and feature detection in most cases (jain1991unsupervised). In the two-dimensional domain, the real Gabor function can be defined as:
@@ 2.2 Squeeze-and-Excitation attention
In a neural network, convolutional blocks are feature extraction layers that fuse spatial and channel-wise information into a set of representations. Despite the good generalisation achieved by these blocks, there is a constant effort to improve their performance on specific tasks, adapting them for target applications. The Squeeze-and-Excitation (SE) (hu2018squeezeandexcitation) block emerges as an alternative solution to this problem, modelling the channels interdependency on the convolutional features. The SE block can be understood as a channel-wise attention mechanism that aggregates and recalibrates the channel-wise feature responses.
@@
With a slight increase in computational cost, the squeeze-and-excitation operation applied to the ImageNet dataset achieved a significant improvement of nearly 25% over previous models (hu2018squeezeandexcitation). When applied to spectral representations, it enhances channel-wise information by capturing nonlinear relationships between multiple two-dimensional features. By combining both channel attention and Gabor convolutional layers within a neural network architecture, it becomes possible to adaptively select two-dimensional frequency bands.
@@ 2.3 Classification tasks in underwater acoustics
In underwater acoustic classification using machine learning frameworks, several combinations of hyper-parameters are found in the literature. They range from audio segment duration and number of classes to the metric calculation process, making direct comparisons of such works difficult and sometimes inappropriate. In addition, some results are often incompatible due to the various manners in which the training and test datasets are divided. Considering this, three classification tasks have been proposed (liu2021underwater): Task 1 consists of first segmenting the original audio into smaller parts, randomising these segments, and dividing them further into training and test subsets; Task 2 utilises the beginning and end of each audio file for training or testing, and then segments it into smaller parts; and Task 3 consists of using each of the entire audios as either a training or test sample, then segmenting it into smaller parts. The tasks gradually increase in difficulty, providing a broader view of the classification problem.
@@ 3.1 Gabor Squeeze-and-Excitation ResNeXt
The initialisation of the Gabor function parameters followed the strategy proposed in (alekseev2019gabornet; meshgini2012face). Considering Equation 1 as the Gabor kernel, the aspect ratio $\gamma$ was set to 1, and the relationship between $\sigma$ and $\omega$, which is the angular frequency ($\omega=2\pi f$), was initialised as $\sigma=\pi/\omega$. The frequencies and orientations of the kernels are initialised using the following equations:
@@
Table 1: The GSE ResNeXt architecture. C=32 denotes the cardinality of the grouped convolutions. The first layer, Gabor Conv1, represents the learnable Gabor convolution. Table inspired by the one provided in xie2017aggregated.

### v2

@@ Abstract
Remotely detecting and classifying underwater acoustic targets is critical for environmental monitoring and defence. However, the complexity of ship-radiated and environmental noise poses significant challenges for accurate signal processing. While recent advancements in machine learning have improved classification accuracy, limited dataset availability and a lack of standardised experimentation hinder generalisation and robustness. This paper introduces GSE ResNeXt, a deep learning architecture integrating learnable Gabor convolutional layers with a ResNeXt backbone enhanced by squeeze-and-excitation attention. The Gabor filters serve as two-dimensional adaptive band-pass filters, extending the feature channel representation. Its combination with channel attention improves training stability and convergence while enhancing the model’s ability to extract discriminative features. The model is evaluated using three training–test split strategies that reflect increasingly complex classification tasks, demonstrating how systematic evaluation design addresses issues such as data leakage, temporal separation, and taxonomy. Results show that GSE ResNeXt consistently outperforms baseline models like Xception, ResNet, and MobileNetV2, in terms of classification performance. Regarding stability and convergence, adding Gabor convolutions to the initial layers of the model reduced training time by up to 62%. During the evaluation of training-testing splits, temporal separation between subsets significantly affected performance, proving more influential than training data volume. These findings suggest that signal processing can enhance model reliability and generalisation under varying environmental conditions, particularly in data-limited underwater acoustic classification. Future developments should focus on mitigating environmental effects on input signals.
@@ 1 Introduction
Remote detection of underwater acoustic targets is a valuable and fundamental asset for environmental protection and defence, as its applications include surveillance, dredging, hydrography, marine life research, and fisheries protection, to name a few [4]. Given the sensitive nature of these applications, the correct detection, processing, and classification of such signals is essential. In the defence field, for instance, incorrect classification could lead to false alarms or even to the non-identification of a potential threat. In environmental protection contexts, misclassifications can directly affect species population estimates, possibly leading to inappropriate management strategies [6].
@@
Recently, a significant increase in anthropogenic activities conducted in the ocean has elevated the levels of underwater acoustic noise [39, 20], driving new studies to focus on the control of maritime traffic by detecting and classifying ship types by their radiated noise. The complexity of such signals, on the other hand, poses a challenge to their effective management in certain situations. Noise recorded by passive sonars is usually composed of environmental sounds and ship-radiated noise [11]. Environmental sounds include sea surface noise, such as breaking waves and wind, and intermittent biological sounds, which depend on factors such as location, weather conditions, and seasonality [55]. Ship-emitted noise, on the other hand, consists of machinery and propeller noises, which appear as constant lower components in the frequency spectra, and hydrodynamic noise, which is a broadband component that shifts between mid and high frequencies according to the ship’s speed [7]. Additionally, multi-path sound reflection caused by interaction with different media, such as the sea surface, can affect classification accuracy within recordings [49, 33].
@@
With the evolution of machine learning methods, their application in underwater target classification has been increasing. With a focus on feature extraction and classification of ship types, recent research has shown an interest in improving the machine learning architecture to extract relevant information from complex input data. A subset of works uses autoencoders to improve the feature extraction process [30, 28, 23], reducing the need for manual hyperparameters tuning. These architectures take advantage of the data-driven training process, which automatically extracts features from the input data, minimising a given loss function. However, the acquisition of such signals is sometimes sensitive and complex, leading to limited available data and, therefore, limited features. Another common strategy applied for detection and classification is the use of recurrent neural network blocks to extract the temporal dependency of the acoustic signal [29, 2, 18]. This strategy aims to extract features based on the temporal components of the audio; however, as the ship noise is almost constant over short periods of time, the improvements by focusing on short-time dependencies are minimal.
@@
Notably, the most common approach to detecting and classifying underwater acoustic signals involves an initial pre-processing stage, where the raw waveform is transformed into a spectrogram. The resulting two-dimensional representation allows the signal to be classified using algorithms originally developed for computer vision tasks. Residual neural networks (ResNet) [19] serve as the baseline for several pipelines in underwater acoustic target classification [21, 58, 38, 12]. They introduce identity shortcut connections between convolutional layers, enabling the network to bypass one or more layers by directly passing the output of earlier layers to later ones. These connections extend the model’s capacity by allowing the exploration of alternative paths during inference, thereby mitigating the vanishing gradient problem [45]. Subsequent improvements introduced the concept of cardinality [51], which enhances the model’s ability to explore multiple learning paths through grouped convolution blocks within residual layers. The resulting model, ResNeXt, often achieves higher classification accuracy with only a modest increase in computational cost. In underwater acoustics, the application of ResNets has proven highly beneficial for improving classification accuracy, as they enable deeper architectures and facilitate the exploration of multiple inference paths [21, 58, 38, 12]. In a similar fashion, multi-branch classification approaches are also frequently employed [53, 52, 46, 43, 37], further enhancing the multi-path capability of the pipelines. The successful application of these strategies highlights the feasibility of employing computer vision techniques for classification tasks based on spectrograms.
@@
Attention mechanisms have been incorporated into classification pipelines, allowing models to dynamically assign importance to input elements and focus on contextually relevant information when producing outputs [3]. These mechanisms have been further emphasised with the advent of transformer models [44], which rely solely on attention mechanisms and dispense with convolutions. Over the years, transformer models have been optimised for mobile applications [34], and variants have also been applied to spectrogram classification [16], highlighting their versatility in contexts outside the field of computer vision. Despite the impressive results obtained in vision applications using transformers, the combination of attention mechanisms and convolutional neural networks is still being explored, yielding superior results in several tasks [17, 25, 41]. In vision tasks, attention is typically applied along the spatial dimension, assigning adaptive importance factors across the image [61, 13], or along the channel dimension, prioritising different feature channels within the neural network [48, 22]. Subsequent approaches also focus on combining both strategies in a spatial–channel framework [50, 8], which aims to enhance the model’s capability to represent and extract features from the inputs. For spectrograms, as the spatial dimension has different scales along the vertical and horizontal axes, some applications separate the time and frequency dimensions when applying spatial attention methods [56]. In underwater acoustic classification, attention mechanisms have proven effective in improving accuracy by assigning different levels of importance to features at local and global scales [60, 57, 54, 56, 47]. While these techniques continue to advance classification performance, their effectiveness is highly dependent on the type of spectral representation used [12]. Processing these representations to address the complex behaviour of underwater acoustic signals offers an attractive alternative to improvements in neural networks alone.
@@
Gabor filters, which are band-pass filters well suited to local frequency and orientation analysis [10], have the potential to enhance feature extraction in convolutional neural networks. Their application in underwater acoustic tasks has also been explored in recent years. In a study of whale vocalisations, Gabor filtering was applied to enhance distributed acoustic sensing (DAS) data, improving the signal-to-noise ratios of the inputs [15]. While it supports the broader applicability of these filters in underwater environments, DAS data differs fundamentally from conventional hydrophone data, which was not addressed by this work. In a hydrophone-based classification problem, Gabor filters were applied directly to raw audio as a preprocessing step, generating a single-channel spectral representation as input [14]. This preprocessing stage dynamically emphasised different frequency components, improving classification performance. However, the resulting spectral input was used without further processing, leaving the task of feature extraction entirely to the model. The application of these filters has also been studied in the two-dimensional domain, where they were applied to spectrogram inputs followed by a low-pass filter [38]. Despite achieving a 3.5% improvement in accuracy, this approach focused solely on enhancing the initial representations, without further processing the resulting frequency bands. These results support the use of Gabor filters to improve generalisation, but leave open the question of how the different frequency bands can be more effectively exploited within the neural network architecture. Furthermore, most previous works consider only a single classification task, which limits the generalisation of their findings.
@@
This paper presents a method for underwater acoustic classification that optimises the feature extraction layer, providing an experimentation strategy that addresses three different scenarios and considers the variability of the training dataset. The Gabor Squeeze-and-Excitation ResNeXt model (GSE ResNeXt) introduces the usage of Gabor convolutional filters in the initial layers of the neural network combined with residual layers and channel attention mechanisms in the models’ backbone, showing to be effective at improving the training stability and classification performance across different classification tasks. Additionally, the use of Gabor filters facilitated the convergence of the model, representing a substantial enhancement for future applications that utilise limited number of training examples. In summary, the contributions of this paper are as follows:
@@
    The integration of learnable Gabor convolutional layers with squeeze-and-excitation channel attention, serving as a two-dimensional adaptive band-pass front-end followed by frequency-selective channel weighting. This design stabilises optimisation and enhances discriminative feature learning, thereby improving classification performance.
@@
    The application of learnable Gabor convolutions in the initial layer accelerates convergence, delivering a smoother validation trajectory and achieving steady state earlier without compromising final performance.
@@
    A systematic evaluation of model generalisation across three training–testing split strategies of increasing complexity. This highlights the impact of temporal separation, data leakage and class taxonomy on classification performance. It also establishes a reproducible framework for benchmarking underwater acoustic classifiers.
@@ 2.1 Two-dimensional Gabor convolution
Traditional convolutional blocks perform the convolution of a learnable two-dimensional kernel on an input image, occasionally adding an additional bias term depending on the application [27]. When visualising the convergence of convolutional networks, it was observed that the convolutional kernels in the initial layers of models trained for image classification converged to Gabor-like filters and colour blob representations [26, 59]. These findings suggest that the training process could be enhanced by constraining the spatial kernels to Gabor functions, which possess significantly fewer parameters to optimise and are capable of extracting spatial textures from visual inputs. Building on these conclusions, although several studies have employed Gabor functions to pre-process and extract features from images, the integration of a Gabor filter layer with learnable weights into the back-propagation algorithm was only proposed years later [1]. Applying this approach to image classification tasks demonstrated improved robustness, as the visual kernels were less sensitive to noisy information, while also reducing training complexity.
A Gabor function can be understood as a complex sinusoid modulated by a Gaussian envelope. While the complex sinusoid captures the frequency and phase information of the features, the Gaussian component localises them in space. In image processing, the imaginary part of the sinusoid is often omitted to simplify the calculation process, since it is sufficient for texture and feature detection in most cases [24]. In the two-dimensional domain, the real Gabor function can be defined as:
@@ 2.2 Squeeze-and-Excitation attention
In a neural network, convolutional blocks are feature extraction layers that fuse spatial and channel-wise information into a set of representations. Despite the good generalisation achieved by these blocks, there is a constant effort to improve their performance on specific tasks, adapting them for target applications. The Squeeze-and-Excitation (SE) [22] block emerges as an alternative solution to this problem, modelling the channels interdependency on the convolutional features. The SE block can be understood as a channel-wise attention mechanism that aggregates and recalibrates the channel-wise feature responses.
@@
With a slight increase in computational cost, the squeeze-and-excitation operation applied to the ImageNet dataset achieved a significant improvement of nearly 25% over previous models [22]. When applied to spectral representations, it enhances channel-wise information by capturing nonlinear relationships between multiple two-dimensional features. By combining both channel attention and Gabor convolutional layers within a neural network architecture, it becomes possible to adaptively select two-dimensional frequency bands.
@@ 2.3 Classification tasks in underwater acoustics
In underwater acoustic classification using machine learning frameworks, several combinations of hyper-parameters are found in the literature. They range from audio segment duration and number of classes to the metric calculation process, making direct comparisons of such works difficult and sometimes inappropriate. In addition, some results are often incompatible due to the various manners in which the training and test datasets are divided. Considering this, three classification tasks have been proposed [29]: Task 1 consists of first segmenting the original audio into smaller parts, randomising these segments, and dividing them further into training and test subsets; Task 2 utilises the beginning and end of each audio file for training or testing, and then segments it into smaller parts; and Task 3 consists of using each of the entire audios as either a training or test sample, then segmenting it into smaller parts. The tasks gradually increase in difficulty, providing a broader view of the classification problem.
@@ 3.1 Gabor Squeeze-and-Excitation ResNeXt
During the training procedure, the initialisation of the Gabor function parameters followed the strategy proposed in [1, 35]. Considering Equation 1 as the Gabor kernel, the aspect ratio $\gamma$ was set to 1, and the relationship between $\sigma$ and $\omega$, which is the angular frequency ($\omega=2\pi f$), was initialised as $\sigma=\pi/\omega$. The frequencies and orientations of the kernels are initialised using the following equations:
@@
Table 1: The GSE ResNeXt architecture. C=32 denotes the cardinality of the grouped convolutions. The first layer, Gabor Conv1, represents the learnable Gabor convolution. Table inspired by the one provided in Xie et al. [51].